A car with mass 1 500 kg moves with costant velocity of 36 ms-1. The driver sees a group of cows in front and he immediately steps on the brake pedal and manages to stop the car in 6 seconds. The distance of the cows from the car when the driver spotted them was 160 m. How far are the cows from the car when the car stops.


1
Expert's answer
2021-06-17T15:07:44-0400

Let "d=160\ m" is the distance from the car to the cows when the driver was spotted them, "d_1" is the stopping distance of the car and "d_2" is the distance from the car to the cows when the car stops. Let's first find the deceleration of the car from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{0-36\ \dfrac{m}{s}}{6\ s}=-6\ \dfrac{m}{s^2}."

Then, we can find the stopping distance of the car from another kinematic equation:


"v^2=v_0^2+2ad_1,""d_1=\dfrac{v^2-v_0^2}{2a}=\dfrac{0-(36\ \dfrac{m}{s})^2}{2\cdot(-6\ \dfrac{m}{s^2})}=108\ m."

Finally, we can find the distance from the car to the cows when the car stops:


"d=d_1+d_2,""d_2=d-d_1=160\ m-108\ m=52\ m."